# Two-sample Comparison through Additive Tree Models for Density Ratios

### 11 Mar 2026 v3

@@ 1 Introduction
Beyond the stand-alone two-sample comparison problem which is our focus herein, DRE has also been applied as a component in various learning methods in which two-sample comparison arises as a sub-problem directly or indirectly. A far-from-exhaustive list includes the estimation of the mutual information (Suzuki et al., 2009a, b), conditional density estimation (Sugiyama et al., 2010), covariate shift correction (Shimodaira, 2000; Sugiyama et al., 2007a; Reddi et al., 2015; Stojanov et al., 2019), outlier detection (Hido et al., 2011; Kato and Teshima, 2021), causal inference (Uehara et al., 2020), and change point detection (Kawahara and Sugiyama, 2009; Liu et al., 2013; du Plessis et al., 2015). Additionally, estimating density ratios has been utilized in Markov Chain Monte Carlo (MCMC) for intractable likelihoods (Kaji and Ročková, 2023).
@@
In contrast to this popular perspective, we hold that DRE should not and does not need to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can instead be directly learned from the latter. To establish the direct route, we propose a new loss function called “balancing loss”, which has a functional form similar to the exponential loss for the classification AdaBoost (Friedman et al., 2000) but is minimized by the true density function. As we will show, the resulting estimates are often substantially more accurate than what the density-ratio trick produces, especially when the sample size is unbalanced, that is, when the setting is more challenging for classifiers that aim to maximize the total successful classification rate. We demonstrate the limitation of adopting the classification AdaBoost in the experiments.
In terms of computational strategies, we introduce two boosting algorithms for estimating the additive tree model on density ratios. We use an additive tree ensemble as the function class for approximating the density ratio, by adding tree-based weak learners. The first algorithm is based on forward-stagewise fitting (Hastie et al., 2009) and it operationally resembles AdaBoost. In each iteration, the observations are weighted according to the discrepancies between the current estimate and the actual responses, so that in each step we can fit a new weak learner that focuses on observations with large discrepancies. This algorithm sequentially and greedily minimizes the balancing loss through fitting a single-tree piecewise constant approximation to the density ratio of the weighted samples. The second algorithm follows the general paradigm of gradient boosting (Friedman, 2001).
@@
The remainder of the paper is organized as follows. In Section 2 we start by introducing the balancing loss for learning the density ratio and highlight its connection to the variational form of the squared Hellinger distance and to the exponential loss for binary classification. We then introduce an additive tree model for the (log) density ratio along with two boosting algorithms for fitting this model based on forward stage-wise fitting and gradient boosting, respectively. In Section 3, we introduce a generalized Bayesian inference framework that uses the balancing loss as a (log) pseudo-likelihood and show that one can accomplish Bayesian backfitting sampling on the model using standard algorithms for BART by exploiting the analogy of the pseudo-likelihood to an exponential family. In Section 4, we evaluate the performance of the proposed methods in comparison with the existing approaches in a few simulation settings, followed by a case study that assesses and compares multiple generative models for microbiome compositions in Section 5. We give concluding remarks in Section 6.
@@ 2.1 The Balancing Loss for Density Ratio Learning
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $(r^{*})^{-1/2}$ and $(r^{*})^{1/2}$, their probability over any measurable set $B$ becomes equal. In this sense, we say that the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio, $\sqrt{p/q}$. We then consider the following balancing loss function:
@@ 2.2 Connections to Existing Concepts in Statistical Learning
The first perspective is related to a common technique widely adopted in the machine learning community for estimating density ratios, often referred to as the “density-ratio trick”, which involves turning the problem into a supervised, binary classification task on the two group labels (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. As we discuss below, the balancing loss arises naturally as a variant of the exponential loss, which is used for classification with AdaBoost (Friedman et al., 2000), but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick against scenarios such as when the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space. These will also be validated experimentally later in our numerical examples where a direct comparison with the density ratio trick will be provided.
@@
$\frac{p(x)}{q(x)}=\frac{\mathrm{Pr}(y=-1)}{\mathrm{Pr}(y=1)}\frac{\mathrm{Pr}(y=-1\mid x)}{\mathrm{Pr}(y=1\mid x)},$
@@
The issues with the density-ratio trick caused by different sample sizes could be addressed by applying different weights to the two groups so that the two groups have equal weights in total. Similar ideas of re-weighting the minor/major groups are proposed in the context of classification for unbalanced data, as, for example, the cost-sensitive boosting (Sun et al., 2007). To the best of our knowledge, however, neither its application to density ratio estimation nor, more importantly, its extension for Bayesian inference is discussed in the literature. Moreover, we need to mention that an alternative calibration method of the estimated classification probability specialized for density ratio estimation is proposed in Cranmer et al. (2015), but the performance of AdaBoost is not substantially improved as we see in Section 4.
@@ 2.3 Additive Tree Models and Boosting
    For each $A\in T_{k}$,
    $\exp(\beta_{k}(A))=\sqrt{\frac{\int_{A}pw^{-1}_{k-1}\,d\mu}{\int_{A}qw_{k-1}\,d\mu}}.$
@@
We can also introduce a gradient boosting algorithm (Friedman, 2001). To this end, note that the negative gradients of the empirical loss with respect to the additive estimand $F$ are as follows:
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=-\frac{1}{n_{0}}\exp(-F(x_{i}^{(0)}))=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)}),\ -\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}\exp(F(x_{i}^{(1)}))=-\frac{1}{n_{1}}w(x_{i}^{(1)}),$
which are the pseudo-residuals. One can then follow the recipe provided in (Friedman, 2001). Specifically, in each iteration of the algorithm we fit a regression tree to the pseudo-residuals by minimizing the squared error loss, that is, minimizing the variances of the gradients in the leaf nodes. We again adopt a coarse-to-fine greedy search and for each node $A$, we select the splitting rule with child nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, we obtain the optimal values of the node parameters, which are already given by Proposition 1.
@@
As for supervised problems, the success of tree boosting under the two algorithms above requires each base learner to be weak in the sense that it extracts only a small amount of structure in the density ratio function. Similar to supervised boosting, we can regularize the contribution of each learner to the estimate by introducing a learning rate denoted by $\nu\in(0,1]$. If the current estimate is $w_{k-1}$ and the optimal learner obtained in the $k$th step is $f_{k}$, we can combine them as follows:
@@
3 Generalized Bayesian Inference on Additive Trees for Density Ratios
Tree boosting only offers point estimates for the density ratio function, without providing any associated uncertainty. To achieve uncertainty quantification for the density ratio, in this section, we introduce a generalized Bayesian strategy (Bissiri et al., 2016) to learning the density ratio.
@@ 3.1 Balancing Loss as a Pseudo-likelihood
$L_{n,\tau}(w)=\frac{\exp(-n_{\mathrm{min}}\tau l_{n}(w))}{\tau^{2n_{\mathrm{min}}}}$
where $n_{\min}$ is the sample size of the smaller group $\min(n_{0},n_{1})$, and $\tau>0$ is a “temperature” parameter. The term in the denominator $\tau^{2n_{\mathrm{min}}}$ is added to provide a clearer interpretation of the pseudo-likelihood as a weighted likelihood under exponential distributions, as follows:
@@
where $\zeta_{0}=n_{\mathrm{min}}/n_{0}$ and $\zeta_{1}=n_{\mathrm{min}}/n_{1}$ are the weights taking values in $(0,1]$. If, for example, the first group has a smaller sample size with $n_{0}<n_{1}$, then the observations from the second group are down-weighted to keep the information from the two groups balanced. The effective sample size is controlled not by the total sample size but by $n_{\mathrm{min}}$, and this is natural since the amount of information about the two-sample difference brought by the data crucially depends on the size of the smaller group.
@@
where $\pi_{0}(w)$ is the prior of the balancing function $w$. Its specification is detailed in the next section.
@@ 3.2.1 Updating The Leaf Node Parameters
where $\gamma_{k}(A)=\exp(\beta_{k}(A))$. Noting that this pseudo-likelihood resembles an exponential family kernel for $\gamma_{k}(A)$, we adopt an inverse-Gaussian leaf prior (Murray, 2021) for $\gamma_{k}(A)$, which is conjugate to this kernel. Specifically, the prior $\mathrm{InverseGauss}(\mu_{k}(A),\lambda_{k}(A))$ on $\gamma_{k}(A)$ takes the following form:
@@ 3.3 Tuning the Temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ relative to the prior and significantly influences the resulting posteriors, including posterior variances and credible intervals. There is a growing body of literature on strategies for selecting the temperature parameter empirically. Some proposed methods include information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). However, it is not straightforward to adapt these methods to the current context, where the task is to learn of an unknown, nonparametrically specified function in an unsupervised setting and the computation cost of the MCMC is relatively expensive.
@@ 3.4 Empirical Validation of Uncertainty Calibration
The posterior distributions under the balanced and unbalanced sample sizes are provided in Figure 1, where the posterior means and the credible intervals are evaluated pointwise. We can see that the posterior distributions are distributed around the true log-densities in a region with an adequate number of observations (i.e., the region between -2 and 2). Moreover, the inverse temperature $\tau^{-1}$, with which we estimate the Bhattacharyya coefficient, is distributed around the true value. We note that more information about the two-sample difference is provided under the balanced sample size as is represented by the weighted likelihood expression, and it is reflected by the different widths of the credible intervals in the posterior distributions. Figure 1 also shows the calibration of the credible intervals obtained by our Bayesian model in terms of nominal coverage rates based on 50 simulated data sets. (For each simulation, the coverage rate is calculated based on the ratio of the observations included in the corresponding credible intervals.) It implies that the posterior uncertainty and the nominal coverage rate tend to agree more under a larger sample size, as in the case of $(n_{0},n_{1})=(2500,2500)$. It is noteworthy that the results under $(500,500)$ and $(500,4500)$ are similar, and again it is consistent with the specification of the loss-based likelihood, in which the amount of information is controlled by the size of the smaller group ($n_{0}=500$ in this example).
@@
Figure 1: An evaluation of posterior distributions under the 1D scenarios with the balanced/unbalanced sample size. Each left plot compares the true log-densities (gray, dotted) and the posteriors (mean: black, solid, 95% interval: blue, dashed) evaluated pointwise. Each middle plot shows the posterior distributions of the inverse temperature $\tau^{-1}$ and the true Bhattacharyya coefficient. Each right plot visualizes the calibration plot comparing the posterior uncertainty and the nominal coverage rate (namely, the ratio of observations included in the pointwise credible intervals) based on 50 simulations and their averages.
@@ 4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases, providing a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost for binary classification, as it uses the same class of additive tree ensembles to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms is essentially due to the different loss functions employed. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition to the density-ratio trick, we evaluate the calibration method proposed in Cranmer et al. (2015) using AdaBoost as a discriminator. We also compare our methods with two popular kernel-based density ratio estimators, namely KLIEP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019).
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed set to 1000. In both types of the boosting algorithms, we introduce 31 equally spaced cut points in each dimension. For a fair comparison, AdaBoost is implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC sampler is initialized with trees containing only root nodes, and the size of the burn-in and the posterior sample is set to 2000 and 1000, respectively. Further details on the implementation are provided in Supplementary Materials C.
@@
and for each setting we calculate the average of the error defined above based on 50 simulated data sets.

### 22 Apr 2026 v4

@@ 1 Introduction
Beyond the stand-alone two-sample comparison problem which is our focus herein, DRE has also been applied as a component in various learning methods in which two-sample comparison arises as a sub-problem directly or indirectly. A far-from-exhaustive list includes covariate shift correction (Shimodaira, 2000; Sugiyama et al., 2007a; Reddi et al., 2015; Stojanov et al., 2019), outlier detection (Hido et al., 2011; Kato and Teshima, 2021), causal inference (Uehara et al., 2020), and change point detection (Kawahara and Sugiyama, 2009; Liu et al., 2013; du Plessis et al., 2015). Additionally, estimating density ratios has been utilized in Markov Chain Monte Carlo (MCMC) for intractable likelihoods (Kaji and Ročková, 2023).
@@
In contrast to this popular perspective, we hold that DRE should not and does not need to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can instead be directly learned from the latter. To establish the direct route, we propose a new loss function called “balancing loss”, which has a functional form similar to the exponential loss used in the classification AdaBoost (Friedman et al., 2000) but is minimized by the true density function. The resulting estimates are often substantially more accurate than what the density-ratio trick produces, especially when the sample size is unbalanced, that is, when the setting is more challenging for classifiers that aim to maximize the total successful classification rate. We demonstrate the limitation of adopting the classification AdaBoost in the experiments.
In terms of computational strategies, we introduce two boosting algorithms for estimating the additive tree model on density ratios. We use an additive tree ensemble as the function class for approximating the density ratio, by adding tree-based weak learners. The first algorithm is based on forward-stagewise fitting (Hastie et al., 2009) and it operationally resembles AdaBoost by minimizing the balancing loss by repeatedly fitting a single-tree weak learner to reweighted observations. The second algorithm follows the general paradigm of gradient boosting (Friedman, 2001).
@@
The remainder of the paper is organized as follows. Section 2 formulates the balancing loss and introduces the boosting algorithms. Section 3 proposes the generalized Bayesian inference based on the balancing loss and the BART model designed for density ratio estimation. Sections 4 and 5 present numerical experiments and an application to the microbiome data, and Section 6 concludes.
@@ 2.1 The Balancing Loss for Density Ratio Learning
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $(r^{*})^{-1/2}$ and $(r^{*})^{1/2}$, their weighted measures over any measurable set $B$ becomes equal. In this sense, we say that the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio, $\sqrt{p/q}$. We then consider the following balancing loss function:
@@ 2.2 Connections to Existing Concepts in Statistical Learning
A first connection is to the “density-ratio trick”, which estimates the density ratio by inverting a binary classifier trained to distinguish the two samples (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. As we discuss below, the balancing loss arises naturally as a variant of the exponential loss, which is used for classification with AdaBoost (Friedman et al., 2000), but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick to scenarios in which the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space.
@@
$\frac{p(x)}{q(x)}=\frac{\mathrm{Pr}(y=-1)}{\mathrm{Pr}(y=1)}\frac{\mathrm{Pr}(y=1\mid x)}{\mathrm{Pr}(y=-1\mid x)},$
@@
The issues with the density-ratio trick caused by different sample sizes could be addressed by applying different weights to the two groups so that the two groups have equal weights in total. Similar ideas of re-weighting the minor/major groups are proposed in the context of classification for unbalanced data, as, for example, the cost-sensitive boosting (Sun et al., 2007). To the best of our knowledge, the literature has not discussed either its application to density ratio estimation or, more importantly, its extension to Bayesian inference.
@@ 2.3 Additive Tree Models and Boosting
    For each $A\in T_{k}$, $\exp(\beta_{k}(A))=\sqrt{\int_{A}pw^{-1}_{k-1}\,d\mu}/\sqrt{\int_{A}qw_{k-1}\,d\mu}$.
@@
We can also introduce a gradient boosting algorithm (Friedman, 2001), noting that the negative gradients of the empirical loss with respect to the additive estimand $F$ are as follows:
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=\frac{1}{n_{0}}\exp(-F(x_{i}^{(0)}))=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)}),\ -\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}\exp(F(x_{i}^{(1)}))=-\frac{1}{n_{1}}w(x_{i}^{(1)}),$
which are the pseudo-residuals. One can then follow the recipe provided in (Friedman, 2001). Specifically, in each iteration of the algorithm we fit a regression tree to the pseudo-residuals by minimizing the squared error loss, that is, minimizing the variances of the gradients in the leaf nodes. We again adopt a coarse-to-fine greedy search and for each node $A$, we select the splitting rule with child nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, we obtain the optimal values of the node parameters given by Proposition 1.
@@
As in supervised problems, the success of tree boosting under the two algorithms above requires each base learner to be regularized to extract only a small amount of structure in the density ratio function. We can regularize the contribution of each learner to the estimate by introducing a learning rate denoted by $\nu\in(0,1]$. If the current estimate is $w_{k-1}$ and the optimal learner obtained in the $k$th step is $f_{k}$, we can combine them as follows:
@@
3 Generalized Bayesian Inference on Additive Trees
Tree boosting only offers point estimates for the density ratio function, without providing any associated uncertainty. To achieve uncertainty quantification, next we introduce a generalized Bayesian strategy (Bissiri et al., 2016) for learning the density ratio.
@@ 3.1 Balancing Loss as a Pseudo-likelihood
$L_{n,\tau}(w)=\tau^{2n_{\mathrm{min}}}\exp(-n_{\mathrm{min}}\tau l_{n}(w))$
where $n_{\min}$ is the sample size of the smaller group $\min(n_{0},n_{1})$, and $\tau>0$ is a “temperature” parameter. The term $\tau^{2n_{\mathrm{min}}}$ is added to provide a clearer interpretation of the pseudo-likelihood as a weighted likelihood under exponential distributions, as follows:
@@
where $\zeta_{0}=n_{\mathrm{min}}/n_{0}$ and $\zeta_{1}=n_{\mathrm{min}}/n_{1}$ are the weights taking values in $(0,1]$. If, for example, the first group has a smaller sample size with $n_{0}<n_{1}$, then the observations from the second group are down-weighted to keep the information from the two groups balanced. The effective sample size is controlled not by the total sample size but by $n_{\mathrm{min}}$, and this is natural since the amount of information about the two-sample difference provided by the data crucially depends on the size of the smaller group.
@@
where $\pi_{0}(w)$ is the prior of the balancing function $w$. Its specification is detailed next.
@@ 3.2.1 Updating The Leaf Node Parameters
where $\gamma_{k}(A)=\exp(\beta_{k}(A))$. Noting that this pseudo-likelihood resembles an exponential family kernel for $\gamma_{k}(A)$, we adopt an inverse-Gaussian leaf prior (Murray, 2021) for $\gamma_{k}(A)$, which is conjugate to this kernel. Specifically, the prior $\mathrm{InverseGauss}(\mu_{k}(A),\lambda_{k}(A))$ on $\gamma_{k}(A)$ is:
@@ 3.3 Tuning the Temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ relative to the prior and significantly influences the resulting posteriors, including posterior variances and credible intervals. There is a growing body of literature on selecting the temperature parameter empirically, such as information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). However, it is not straightforward to adapt these methods to the current context, where the task is to learn an unknown, nonparametrically specified function in an unsupervised setting and the computation cost of the MCMC is relatively expensive.
@@ 3.4 Empirical Validation of Uncertainty Calibration
The posterior distributions under the balanced and unbalanced sample sizes are provided in Figure 1, where the posterior means and the credible intervals are evaluated pointwise. We can see that the posterior distributions are centered around the true log-densities in a region with an adequate number of observations (i.e., the region between -2 and 2). Moreover, the inverse temperature $\tau^{-1}$, with which we estimate the Bhattacharyya coefficient, is distributed around the true value. We note that more information about the two-sample difference is provided under the balanced sample size as is represented by the weighted likelihood expression, and it is reflected by the different widths of the credible intervals in the posterior distributions. Figure 1 also shows the calibration of the credible intervals obtained by our Bayesian model in terms of nominal coverage rates based on 50 simulated data sets. It implies that the posterior uncertainty and the nominal coverage rate tend to agree more under a larger sample size, as in the case of $(n_{0},n_{1})=(2500,2500)$. The cases $(500,500)$ and $(500,4500)$ are similar, and it is consistent with the loss-based likelihood being driven by the smaller group size.
@@
Figure 1: An evaluation of posterior distributions under the 1D scenarios with the balanced/unbalanced sample sizes. Each left plot compares the true log-density ratios (gray, dotted) and the posteriors (mean: black, solid, 95% interval: blue, dashed) evaluated pointwise. Each middle plot shows the posterior distributions of the inverse temperature $\tau^{-1}$ and the true Bhattacharyya coefficient. Each right plot visualizes the calibration plot comparing the posterior uncertainty and the nominal coverage rate (namely, the ratio of observations included in the pointwise credible intervals) based on 50 simulations and their averages.
@@ 4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases, providing a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost for binary classification, as it uses the same class of additive tree ensembles to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms is essentially due to the different loss functions employed. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition to the density-ratio trick, we evaluate the calibration method proposed in Cranmer et al. (2015) using AdaBoost as a discriminator. We also compare our methods with two popular kernel-based density ratio estimators, namely KLIEP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019). For the proposed methods, we assess both types of boosting, the FS and the GB algorithms, along with the Bayesian additive tree model. Their implementation details are provided in Supplementary Materials C.
@@
and for each setting we calculate the average error defined above based on 50 simulated datasets.